Does the force constant of a spring changes if it is cut into parts?

Grade:12th Pass

14 Answers

Prajwal kr
11 years ago

Yes.

Akash Kumar Dutta
11 years ago

no.

it only depends upon the nature of the material.

MUKESH chinna YADAV
11 years ago

force constant does not change

Rohan Das
11 years ago

we know F=-Kx

so force is proportional to length but k(spring constant) depends upon nature of spring

since the spring remains the same k will remain same for both the parts.

MUKESH chinna YADAV
11 years ago

force constant does not changes

Chetan Mandayam Nayakar
11 years ago

Yes. The force constant is inversely proportional to length.

gauhar singh
11 years ago
yes
venkatesh vadde
11 years ago

No, it will not change.

JAYANT ARORA
11 years ago

 We know that F=k.x , therefore k=F/x ..... now if we cut the spring into 2 parts .... F will increase as x is halfed ....

kasireddy siva
11 years ago

the force constant of a spring of does not change.Because,it has a "inherent property" along with it.Due this the force constant doesn''t change.

Abhijeet Vyas
11 years ago

Not at all

Mohit Gupta
11 years ago

Dear Anubhab,

force constant of a spring do change when you cut it into parts. Spring constant of a spring (k) depends upon three factors 

- length of spring

- radius of spring

- material of spring

it is inversely proportional to the length of spring so if you cut the spring into parts the spring constant will increase.

Hope this helps 

Akash Kumar Dutta
11 years ago

i am sorry.the ans is that yes it DOES CHANGES.

take an example.

when two springs of spring constant k are connected end to end then resultant

spring constant is k/2.

HENCE,it can be thought as one single spring with spring constant k/2.

and when it is cut to two parts the individual spring constant is k.

the same way when a spring of spring constant k is cut into n parts,the new spring constants are nk of the individual springs.

and when the same force is applied to the springs.then x=F/nk(less than original extensin=F/k by a factor of 1/n).

this explains that smaller springs extend less than the original springs,when applied by the same force.

HOPE I HELPED YOU.

pvn s
11 years ago

when spring is cut into two parts, then the force constant (k previusly) becomes 2k for each half

 approve this anwer, if it helpsWink